# Fine-Tuning Code Language Models to Detect Cross-Language Bugs

### v1

@@ Abstract.
With the rapid development of software engineering technologies, multilingual programming, i.e., developing a program using multiple programming languages (PLs) collaboratively, has become increasingly popular. This practice is often adopted to reuse existing code, to leverage advantages of specific PLs, to satisfy various software quality needs, and to enhance software development efficiency. However, it also introduces cross-language bugs (CLBs), bugs caused by interactions between multiple PLs. Most existing bug detection tools and techniques are primarily designed for a single PL and are less effective at detecting CLBs. Given the significant advancements of Pre-trained Language Models (PLMs) in the code domain, this work investigates the performance of pre-trained Code Language Models (CodeLMs) in CLB prediction tasks. To this end, we designed and developed a cross-language code identification tool (CLCFinder), and used it to collect and construct a CLB dataset, which involves combinations of three PLs (i.e., Python-C/C++, Java-C/C++, and Python-Java) with nine types of interaction mechanisms. We fine-tuned 13 popular CodeLMs in different sizes on this dataset, in which 80% of the data was used for fine-tuning, 10% for validation, and 10% for testing. We then evaluated their performance in CLB detection. Additionally, we compared the fine-tuning performance of five top-performing models on current state-of-the-art (SOTA) single-language bug datasets to assess their application in CLB detection. Furthermore, we evaluated factors affecting the performance of fine-tuned CodeLMs in CLB detection, including the size of the fine-tuning dataset and the length of token sequence. Finally, we explored the impact of code comments on the performance of fine-tuned CodeLMs. The experimental results are that: First, all 13 CodeLMs performed poorly in CLB detection before fine-tuning, but exhibited varying degrees of performance improvement after fine-tuning. In particular, after fine-tuning, the UniXcoder-base model achieved the best performance, with an F1 score of 0.7407. Notably, small CodeLMs tended to perform better than large CodeLMs, while large CodeLMs showed limited improvement. Second, the fine-tuned CodeLMs on single-language bug datasets performed poorly in CLB detection, further demonstrating the significant feature differences between CLBs and single-language bugs. Third, enlarging the fine-tuning dataset size significantly enhanced the model performance in CLB detection, but increasing token sequence length did not necessarily improve the model performance. Finally, regarding the impact of code comments, the performance varied significantly across different CodeLMs. Specifically, some CodeLMs’ performance was improved when fine-tuned with code comments, while others showed degraded performance.
@@ 1. Introduction
As the scale and complexity of software systems continue to grow, multilingual programming has become a critical practice in modern software development. By combining the strengths of different programming languages (PLs), developers can build highly scalable and high-performance software systems more efficiently. However, this multilingual programming paradigm also introduces new challenges - cross-language bugs (CLBs). In a multilingual programming environment, interactions between different PLs are typically facilitated through interfaces, libraries, or middleware (Li et al., 2023a). For example, Python and C/C++ may interact through Cython or ctypes, while Java and C/C++ often interact via the Java Native Interface (JNI). These languages interaction mechanisms offer powerful functionalities, but also introduce significant complexity (Abidi et al., 2021). CLBs typically arise in interfaces or interaction mechanisms between different PLs, and due to their concealed nature and complexity, traditional single-language defect detection methods are often inadequate in effectively addressing CLBs (McCormack et al., 2024). For example, if a Python script calls a C++ library and the memory allocation in the C++ library is not properly released, it may lead to memory leak issues. In a single-language environment, such issues can be relatively easily detected using static code analysis tools. However, in a cross-language scenario, due to the lack of unified tool chains and debugging support, these issues often become difficult to locate and fix. Such problems involve not only the language itself, but also the interface design and runtime behavior between different PLs, making bug detection a more challenging task (Li et al., 2023b).
@@
Existing bug detection research focuses primarily on single-language scenarios. Machine learning-based bug detection methods, such as Support Vector Machines (SVM) and Random Forests (RF), rely on manual feature extraction and often struggle to capture the semantic information of the code (Li et al., 2024c). Although deep learning models, such as Convolutional Neural Networks (CNN)  (Liu et al., 2023), have made significant progress in single-language bug detection, these methods typically overlook the complexity of interactions between multiple PLs and have not been optimized for CLBs (Giray et al., 2023). Despite that recent studies have attempted to explore the possibility of multilingual bug detection, such as the just-in-time multilingual bug prediction technique (Li et al., 2024c), they focused mainly on multilingual bugs but have not effectively addressed bugs arising from cross-language interactions. Moreover, in the cross-language bug dataset constructed by Yang et al. (Yang et al., 2024), the Python code samples are bug-free code, meaning that the trained model cannot detect bugs within Python code.
With the rapid development of Large Language Models (LLMs), Pre-trained Language Models (PLMs) (such as GPT (Radford et al., 2018), BERT (Devlin et al., 2019)) have achieved significant success in the field of natural language processing (NLP). Similarly, pre-trained Code Language Models (CodeLMs) (such as CodeBERT (Feng et al., 2020) and CodeLlama (Roziere et al., 2023)) have excelled in tasks such as code understanding, generation, and completion, enabling a deep understanding of the context of the code and consequently facilitating the detection of potential defects. This opens up the possibility of utilizing CodeLMs for CLB detection. However, how to effectively leverage these CodeLMs for CLB detection remains an open challenge that requires further exploration.
@@ 2.1. Software Bug Detection
Software bug detection is a key task for ensuring software quality and has evolved from traditional static analysis and dynamic analysis to the current approaches leveraging machine learning and CodeLMs (Wang et al., 2024; Li et al., 2024a, c; Mo et al., 2022).
Common static analysis tools include Error Prone from Google (Inc., [n. d.]), Infer from Facebook (Facebook, 2013), and SpotBugs (SpotBugs, 2021), the successor to the widely used FindBugs tool (Ayewah et al., 2008), are examples of popular static bug detection tools. Recent years, researchers have sought to enhance the comprehensiveness and accuracy of bug detection by evaluating and integrating different types of bug detection techniques, such as flow pattern matching (Hovemeyer and Pugh, 2004), abstract interpretation (Cousot and Cousot, 1977), symbolic execution (King, 1976), and program model checking (Clarke et al., 2009).
@@
With the rapid advancement of machine learning and deep learning techniques and their demonstrated effectiveness in various domains, researchers have increasingly explored their application in software bug detection, capitalizing on the growing volume of fault data generated during the software development process (Liu et al., 2023; Deng et al., 2020; Li et al., 2017; Giray et al., 2023). Li et al. proposed a software bug prediction tool based on the RF algorithm, aimed at predicting whether code changes introduce bugs involving multiple PLs in a timely manner (Li et al., 2024c). Wongpheng et al. utilized a multi-layer CNN to extract features from code snippets, achieving higher detection accuracy compared to traditional machine learning methods (Wongpheng and Visutsak, 2020). Similarly, Khleel et al. employed a variant of RNN—Gated Recurrent Unit (GRU)—to capture temporal dependencies in code, combining it with CNNs for feature extraction, thereby improving detection performance (Khleel and Nehéz, 2023).
@@ 2.2. Multilingual Bug and CLB Detection
With the widespread use of multilingual software systems, an increasing number of studies have focused on bug detection and security analysis in multilingual software systems. Research indicates that the interaction mechanisms between PLs in multilingual software systems significantly impact software development, with the bug rate in cross-language dependencies being three times higher and the vulnerability rate being twice that of single-language dependencies (Li et al., 2024d, 2022a). Similarly, in their study of cross-language dependencies, Grichi et al. found that cross-language dependencies are more likely to introduce bug and security vulnerabilities than single-language dependencies (Grichi et al., 2020). This highlights the importance of cross-language software analysis for software security, which has been actively studied. Sungjae Hwang et al. developed JUSTGen, a semi-automated tool, to study the behavior of JVM when handling JNI errors (Hwang et al., 2024, 2021). In addition, a decompilation method has been proposed for statically analyzing JNI programs, which effectively analyzes compiled JNI programs, overcoming the limitations of traditional static analysis tools (Park et al., 2023). The PolyCrusie tool combines symbolic dependency analysis and data flow analysis to effectively identify cross-language security vulnerabilities, achieving significant results in Python-C programs (Li et al., 2022c). xLoc uses a Transformer-based deep learning model to effectively detect and locate CLBs (Yang et al., 2024). Similarly, the MVD framework proposes an innovative approach by covering six major PLs and learning from a dataset containing over 11,000 real-world vulnerabilities, significantly improving vulnerability detection performance (Zhang et al., 2024a).
@@ 2.3. Large Language Models for Bug Detection
The success of LLMs in natural language processing has led to growing efforts to apply them in software bug detection, resulting in a number of effective and well-evaluated approaches. In static analysis and bug detection, Li et al. proposed the LLIft framework, which combines LLMs with traditional static analysis techniques to enhance the detection of UBI (Use Before Initialization) bugs in the Linux kernel (Li et al., 2024b). Similarly, GPTScan integrates LLMs with static analysis for detecting vulnerabilities in smart contracts (Sun et al., 2024). Khare et al. conducted a comprehensive analysis of LLMs’ vulnerability detection capabilities and found that in some cases, LLMs perform similarly to static analysis tools like CodeQL (Khare et al., 2023). Shu et al. evaluated PLMs and LLMs on multi-language vulnerability detection using over 30,000 real patches, demonstrating LLMs’ superior cross-language semantic understanding and high-risk vulnerability identification (Shu et al., 2025).
@@ 2.4. Conclusive Summary
In recent years, CodeLMs have demonstrated outstanding performance in software engineering tasks, including code understanding, code generation, and single-language bug detection (Feng et al., 2020; Guo et al., 2020; Bai et al., 2023; Hui et al., 2024; Guo et al., 2024c). Based on this trend, we attempted to apply CodeLMs to CLB detection tasks. We developed CLCFinder and built a CLB dataset based on it. Using this dataset, we fine-tuned 13 prevailing CodeLMs and conducted a set of experiments, with the detailed experimental design presented in the following section (Section 3).
@@ 3.1. Objective and Research Questions
The goal of this study is to investigate the performance of the prevailing CodeLMs in detecting CLBs before and after fine-tuning, as well as the factors that influence their performance. We first fine-tuned several popular CodeLMs on our CLB dataset and evaluated their effectiveness on the cross-language detection task to validate the applicability and superiority of the dataset for this purpose. We then evaluated the performance of CodeLMs fine-tuned on mainstream single-language bug datasets when applied to cross-language scenarios, to further highlight the distinctiveness and necessity of our CLB dataset compared to existing single-language bug datasets. We subsequently designed a series of experiments to systematically analyze the factors affecting the models’ detection performance, given the varying performance of different CodeLMs in CLB detection and the fact that our dataset exhibits imbalances in the code length distribution while being significantly smaller in size compared to the pre-training datasets of the CodeLMs. Specifically, we examined the impact of the fine-tuning dataset size and the token sequence length on the performance of the fine-tuned CodeLMs in CLB detection. Finally, we further investigated the impact of code comments — representing natural language elements — during fine-tuning, given that CodeLMs are pre-trained on both code and natural language texts, aiming to evaluate the effect of code comments on CLB detection performance.
@@
RQ2: Can CodeLMs fine-tuned on the current SOTA single-language bug datasets detect CLBs? What is their performance? Prior research on bug detection has primarily relied on single-language bug datasets (Guo et al., 2024a; Yang et al., 2024). However, in multilingual codebases, the transferability of such models is uncertain. With this RQ, we investigate whether CodeLMs fine-tuned on widely adopted single-language bug datasets can generalize to the CLB detection task. By comparing their performance with those trained directly on our CLB dataset, we aim to highlight the limitations of single-language training in cross-language contexts.

### v2

@@ Abstract.
With the rapid development of software engineering technologies, multilingual programming, i.e., developing a program using multiple programming languages (PLs) collaboratively, has become increasingly popular. This practice is often adopted to reuse existing code, to leverage advantages of specific PLs, to satisfy various software quality needs, and to enhance software development efficiency. However, it also introduces cross-language bugs (CLBs), bugs caused by interactions between multiple PLs. Most existing bug detection tools and techniques are primarily designed for a single PL and are less effective at detecting CLBs. Given the significant advancements of Pre-trained Language Models (PLMs) in the code domain, this work investigates the performance of pre-trained Code Language Models (CodeLMs) in CLB prediction tasks. To this end, we designed and developed a cross-language code identification tool (CLCFinder), and used it to collect and construct a CLB dataset, which involves combinations of three PLs (i.e., Python-C/C++, Java-C/C++, and Python-Java) with nine types of interaction mechanisms. We fine-tuned 13 popular CodeLMs in different sizes on this dataset, in which 80% of the data was used for fine-tuning, 10% for validation, and 10% for testing. We then evaluated their performance in CLB detection. Additionally, we compared the fine-tuning performance of five top-performing models on current state-of-the-art (SOTA) single-language bug datasets to assess their application in CLB detection. Furthermore, we evaluated factors affecting the performance of fine-tuned CodeLMs in CLB detection, including the size of the fine-tuning dataset and the length of token sequence. Finally, we explored the impact of code comments on the performance of fine-tuned CodeLMs. The experimental results are that: First, all 13 CodeLMs exhibited varying degrees of performance improvement after fine-tuning. In particular, after fine-tuning, the UniXcoder-base model achieved the best performance, with an F1 score of 0.7407. Notably, within the scope of our experimental setup, small CodeLMs tended to achieve better performance than large CodeLMs, while large CodeLMs showed limited improvement. Second, the fine-tuned CodeLMs on single-language bug datasets performed poorly in CLB detection, further demonstrating the significant feature differences between CLBs and single-language bugs. Third, enlarging the fine-tuning dataset size significantly enhanced the model performance in CLB detection, but increasing token sequence length did not necessarily improve the model performance. Finally, regarding the impact of code comments, the performance varied significantly across different CodeLMs. Specifically, some CodeLMs’ performance was improved when fine-tuned with code comments, while others showed degraded performance.
@@ 1. Introduction
As modern software systems continue to grow in scale and complexity, developing software using multiple interacting programming languages (PLs) has become increasingly common in both industry and the open source community (Delorey et al., 2007; Capers, 2010; Li et al., 2021; Mayer and Bauer, 2015; Tomassetti and Torchiano, 2014). By combining the strengths of different PLs, developers can build highly scalable and high performance software systems more efficiently. Since each PL has its own strengths and limitations, combining multiple PLs is often a necessary choice to achieve functionality or performance requirements that cannot be satisfied by a single-language alone (Abidi et al., 2019; Bissyandé et al., 2013; Meyerovich and Rabkin, 2013; Vasilescu et al., 2013; Yang et al., 2023). As a result, multilingual software systems have gradually become the dominant form of modern software development compared to single-language systems (Li et al., 2021; Mayer and Bauer, 2015; Valverde and Solé, 2015).
However, compared to single-language systems, multilingual software introduces significantly higher complexity. In multilingual software systems, interactions between different PLs are typically implemented through interfaces, libraries, or middleware (Li et al., 2023c). For example, Python can interact with C/C++ through Cython or ctypes, while Java typically interacts with C/C++ through the Java Native Interface (JNI). Although these cross-language interaction mechanisms provide powerful functionality, they also introduce substantial system complexity (Abidi et al., 2021). Compared to single-language systems, the complexity of multilingual software does not only come from the simple aggregation of the complexity of individual languages. More importantly, it arises from the interactions between heterogeneous language components (Abidi et al., 2021; Bae et al., 2019; Grichi et al., 2020; Mayer et al., 2017; Yang et al., 2024). These cross-language interactions lead to a distinct class of bugs that caused originate from language interfaces and cross-language information flows, such as control flow and data flow. These bugs are known as cross-language bugs (CLBs) (Lee et al., 2020; Li et al., 2022c, a). Unlike bugs confined within a single-language, CLBs occur at language boundaries. Fundamental differences in execution models, memory management strategies, and type systems across languages make these bugs more hidden and harder to detect (McCormack et al., 2024; Li et al., 2023b). They also tend to lead to more severe consequences, including memory corruption, security vulnerabilities, and system crashes (Yang and Cai, 2025).
@@
Prior empirical studies on real-world CLBs show that these bugs are not random but arise from recurring root causes at language boundaries (Yang and Cai, 2025). Common causes include misuse of cross-language interfaces and violations of implicit API assumptions, mismatches in memory ownership and lifetime management between managed and unmanaged languages, and semantic inconsistencies in data representation and type conversion. In addition, differences in control flow and error-handling mechanisms across languages can lead to unexpected execution behaviors. CLBs are primarily interface- and interaction-centric, indicating that effective CLB analysis should focus on detecting anomalous cross-language interaction patterns rather than defects confined within a single-language.
Multilingual software may also contain bugs that exist only within a single-language component. Such bugs can usually be effectively handled using mature language specific testing and debugging tools (Li et al., 2022a, 2023a). In contrast, CLBs represent a significant gap in modern software testing and bug diagnosis. And existing bug detection research mainly focuses on single-language scenarios. Current static analysis techniques usually rely on language specific intermediate representations and semantic rules, which makes it difficult for them to analyze about cross-language data flow, control flow, and resource management in a unified way and often leads to high false positive rates (Bae et al., 2019; Lee et al., 2020; Park et al., 2023; Youn et al., 2023; Wei et al., 2018). Dynamic analysis techniques can achieve higher precision, but their effectiveness strongly depends on the availability and coverage of test cases. Real world multilingual systems often lack sufficient testing support (Bai et al., 2018; Li et al., 2022c, 2023a; Xue et al., 2018). In addition, generating high quality test cases typically requires substantial manual effort (Hwang et al., 2021). Consequently, many CLBs cannot be effectively detected by existing static or dynamic techniques and may only surface during system execution. For example, an implicit data type mismatch between Python and a C extension might not trigger any compile-time warnings in either language but could lead to silent data corruption that is only observable under specific runtime conditions.
In recent years, bug detection approaches based on deep learning (DL) have attracted increasing attention (Li et al., 2019; Yang et al., 2022, 2024). Unlike traditional static or dynamic analysis techniques, DL based methods are data driven and do not rely on explicitly encoded language rules or a unified intermediate representation. Since they avoid heavy language specific analysis, they are often easier to extend to multiple language combinations. In particular, rapidly advancing pre-trained code language models (CodeLMs), such as CodeBERT (Feng et al., 2020) and CodeLlama (Roziere et al., 2023), have demonstrated strong capabilities in code understanding, code generation, and code completion tasks. These models can learn rich contextual and semantic representations from large scale code corpora. As discussed above, CLBs are often caused by implicit and undocumented assumptions at language boundaries, such as mismatched memory ownership, inconsistent data representations, or incompatible lifecycle management across languages, which makes these recurring interaction patterns difficult to capture using rule-based analyses. This allows them to implicitly capture recurring patterns in cross-language interactions, making them a promising technical foundation for CLB detection. However, how to effectively leverage CodeLMs for CLB detection remains an open research problem.
Although prior studies have begun to explore multilingual bug detection, existing work mainly focuses on bugs that require changes across multiple languages and does not sufficiently address bugs caused by direct cross-language interactions. For example, just-in-time multilingual bug prediction techniques (Li et al., 2024d) primarily target multilingual bugs rather than CLBs. In addition, in the CLB dataset constructed by Yang et al. (Yang et al., 2024), the Python code samples are bug free. As a result, models trained on this dataset cannot detect bugs in Python code, which limits their effectiveness in realistic cross-language scenarios. Based on these observations, it is necessary to conduct a systematic study of CLB detection and to comprehensively evaluate the effectiveness of CodeLM based approaches for this task.
@@ 2.1. Software Bug Detection
Software bug detection is a key task for ensuring software quality and has evolved from traditional static analysis and dynamic analysis to the current approaches leveraging machine learning and CodeLMs (Wang et al., 2024; Li et al., 2024c, d; Mo et al., 2022).
Common static analysis tools include Error Prone from Google (Inc., ), Infer from Facebook (Facebook, 2013), and SpotBugs (SpotBugs, 2021), the successor to the widely used FindBugs tool (Ayewah et al., 2008), are examples of popular static bug detection tools. Recent years, researchers have sought to enhance the comprehensiveness and accuracy of bug detection by evaluating and integrating different types of bug detection techniques, such as flow pattern matching (Hovemeyer and Pugh, 2004), abstract interpretation (Cousot and Cousot, 1977), symbolic execution (King, 1976), and program model checking (Clarke et al., 2009).
@@
With the rapid advancement of machine learning and deep learning techniques and their demonstrated effectiveness in various domains, researchers have increasingly explored their application in software bug detection, capitalizing on the growing volume of fault data generated during the software development process (Liu et al., 2023; Deng et al., 2020; Li et al., 2017; Giray et al., 2023). Li et al. proposed a software bug prediction tool based on the RF algorithm, aimed at predicting whether code changes introduce bugs involving multiple PLs in a timely manner (Li et al., 2024d). Wongpheng et al. utilized a multi-layer CNN to extract features from code snippets, achieving higher detection accuracy compared to traditional machine learning methods (Wongpheng and Visutsak, 2020). Similarly, Khleel et al. employed a variant of RNN—Gated Recurrent Unit (GRU)—to capture temporal dependencies in code, combining it with CNNs for feature extraction, thereby improving detection performance (Khleel and Nehéz, 2023).
@@ 2.2. Multilingual Bug and CLB Detection
With the widespread use of multilingual software systems, an increasing number of studies have focused on bug detection and security analysis in multilingual software systems. Research indicates that the interaction mechanisms between PLs in multilingual software systems significantly impact software development, with the bug rate in cross-language dependencies being three times higher and the vulnerability rate being twice that of single-language dependencies (Li et al., 2024b, 2022a). Similarly, in their study of cross-language dependencies, Grichi et al. found that cross-language dependencies are more likely to introduce bug and security vulnerabilities than single-language dependencies (Grichi et al., 2020). This highlights the importance of cross-language software analysis for software security, which has been actively studied. Sungjae Hwang et al. developed JUSTGen, a semi-automated tool, to study the behavior of JVM when handling JNI errors (Hwang et al., 2024, 2021). In addition, a decompilation method has been proposed for statically analyzing JNI programs, which effectively analyzes compiled JNI programs, overcoming the limitations of traditional static analysis tools (Park et al., 2023). The PolyCrusie tool combines symbolic dependency analysis and data flow analysis to effectively identify cross-language security vulnerabilities, achieving significant results in Python-C programs (Li et al., 2022c). xLoc uses a Transformer-based deep learning model to effectively detect and locate CLBs (Yang et al., 2024). Similarly, the MVD framework proposes an innovative approach by covering six major PLs and learning from a dataset containing over 11,000 real-world vulnerabilities, significantly improving vulnerability detection performance (Zhang et al., 2024).
@@ 2.3. Large Language Models for Bug Detection
The success of LLMs in natural language processing has led to growing efforts to apply them in software bug detection, resulting in a number of effective and well-evaluated approaches. In static analysis and bug detection, Li et al. proposed the LLIft framework, which combines LLMs with traditional static analysis techniques to enhance the detection of UBI (Use Before Initialization) bugs in the Linux kernel (Li et al., 2024a). Similarly, GPTScan integrates LLMs with static analysis for detecting vulnerabilities in smart contracts (Sun et al., 2024). Khare et al. conducted a comprehensive analysis of LLMs’ vulnerability detection capabilities and found that in some cases, LLMs perform similarly to static analysis tools like CodeQL (Khare et al., 2023). Shu et al. evaluated PLMs and LLMs on multi-language vulnerability detection using over 30,000 real patches, demonstrating LLMs’ superior cross-language semantic understanding and high-risk vulnerability identification (Shu et al., 2025).
@@ 2.4. Conclusive Summary
In recent years, CodeLMs have demonstrated outstanding performance in software engineering tasks, including code understanding, code generation, and single-language bug detection (Feng et al., 2020; Guo et al., 2020; Bai et al., 2023; Hui et al., 2024; Guo et al., 2024a). Based on this trend, we attempted to apply CodeLMs to CLB detection tasks. We developed CLCFinder and built a CLB dataset based on it. Using this dataset, we fine-tuned 13 prevailing CodeLMs and conducted a set of experiments, with the detailed experimental design presented in the following section (Section 3).
@@ 3.1. Objective and Research Questions
The goal of this study is to investigate the performance of the prevailing CodeLMs in detecting CLBs after fine-tuning, as well as the factors that influence their performance. We first fine-tuned several popular CodeLMs on our CLB dataset and evaluated their effectiveness on the cross-language detection task to validate the applicability and superiority of the dataset for this purpose. We then evaluated the performance of CodeLMs fine-tuned on mainstream single-language bug datasets when applied to cross-language scenarios, to further highlight the distinctiveness and necessity of our CLB dataset compared to existing single-language bug datasets. We subsequently designed a series of experiments to systematically analyze the factors affecting the models’ detection performance, given the varying performance of different CodeLMs in CLB detection and the fact that our dataset exhibits imbalances in the code length distribution while being significantly smaller in size compared to the pre-training datasets of the CodeLMs. Specifically, we examined the impact of the fine-tuning dataset size and the token sequence length on the performance of the fine-tuned CodeLMs in CLB detection. Finally, we further investigated the impact of code comments — representing natural language elements — during fine-tuning, given that CodeLMs are pre-trained on both code and natural language texts, aiming to evaluate the effect of code comments on CLB detection performance.
@@
RQ2: Can CodeLMs fine-tuned on the current SOTA single-language bug datasets detect CLBs? What is their performance? Prior research on bug detection has primarily relied on single-language bug datasets (Guo et al., 2024c; Yang et al., 2024). However, in multilingual codebases, the transferability of such models is uncertain. With this RQ, we investigate whether CodeLMs fine-tuned on widely adopted single-language bug datasets can generalize to the CLB detection task. By comparing their performance with those trained directly on our CLB dataset, we aim to highlight the limitations of single-language training in cross-language contexts.